A sample containing 1.90 moles of Ne gas has an initial volume of 7.50 L . What is the final volume of the gas, in liters, when each of the following changes occurs in the quantity of the gas at constant pressure and temperature?

Part A A leak allows one-half of the Ne atoms to escape.

_____L

Part B A sample of 4.00 moles of Ne is added to the 1.90 moles of Ne gas in the container.

____L

Part C A sample of 25.0 g of Ne is added to the 1.90 moles of Ne gas in the container.

____L
Answer #1

initial = 1.90 moles

volume = 7.50 L

volume is directly proportianl to moles

part A)

V1 / V2 = n2 / n2

7.50 /V2 = 1.90 / 0.95

V2 = 3.75

volume = 3.75

part B)

total moles = 4 + 1.90 = 5.90

1.90 / 5.90 = 7.50 / V2

V2 = 23.3 L

volume = 23.3 L

part C)

mass of Ne = 25

moles = 25 / 20 = 1.25

total moles = 3.15

1.90 / 3.15 = 7.5 / V2

V2 = 12.4 L

volume = 12.4 L